# From Landau two-fluid model to de Sitter Universe

### v9

@@ VI.6 Entropy of black hole singularity
De Sitter thermodynamics obeys a modified Gibbs-Duhem relation in the equation (18), which includes the gravitational degrees of freedom: the gravitational coupling $K$ (the gravitational analog of the chemical potential) and the scalar curvature ${\cal R}$. Let us assume that a similar relation applies to the local thermodynamics of black holes. In case of the black hole this relation has the following form:
@@
The same singular entropy density in Eq.(52) is obtained for the charged Reissner-Nordström (RN) black hole. According to Ref.Volovik2022G the entropy and the Bekenstein-Hawking temperature of the RN black hole are the same as for the electrically neutral black hole with the same mass. The extra term is to be added to the right hand side of Eq.(50), $-\Phi q({\bf r})$, which comes from charge density $q({\bf r})=Q\delta({\bf r})$. However this term is zero, because the corresponding electric analogue of the chemical potential (the electrostatic potential) is zero, $\Phi=0$.
So, if our assumption is correct, then it follows from Eqs. (52) and (53) that all the entropy of the black hole, $S_{\rm BH}=A/4G$, is concentrated in the black hole singularity. This is quite natural from the condensed matter point of view: a singularity is a compact physical object with an enormous energy density. The condensed matter analogue of such object is represented by the core of a vortex in a superfluid liquid. The vortex core contains the singularity in the superfluid velocity, ×$\nabla\times{\bf v}_{s}=\kappa\hat{\bf z}\delta({\bf r}_{\perp})$, where $\kappa$ is the quantum of circulation of the superfluid velocity. In the chiral Weyl superfluids, the vortex core has a large density of states, which is caused by the flat band in the spectrum of electrons confined in the singularity.Volovik2011
It would be interesting to apply this approach to the Kerr black hole, and to the super-critical Reissner-Nordström black hole with naked singularity.Tumulka2025
@@ VII.1 Schwinger vs Unruh
The rate of the Schwinger pair creationSchwinger1951 ; Schwinger1953 of particles with mass $M$ and charges ±$\pm q$ in electric field ${\cal E}$ per unit volume per unit time is given by:
@@ VII.2 Double Unruh temperature from Schwinger+Unruh
To connect the Schwinger pair production with the Unruh effect, let us consider the following combined process in which the particles charges ±$\pm q$ with masses $M+m$ are created. At first step the particles with charges ±$\pm q$ with masses $M$ are created, at the second step these charged particles are accelerated by electric field. These particles play the role of the detectors which energy levels are excited from the levels with energy $M$ to the levels with energy $M+m$ due to the Unruh effect. Since these processes occur in unison, the probability of the combined process $\Gamma^{\rm Schw}(M+m)$ is equal to the product of partial probabilities:
@@ VIII.1 Vacuum energy in condensed matter
In condensed matter systems, such as superfluid 4He and superfluid phases of 3He, we know physics both at the ultraviolet (atomic) scale and in the infrared (hydrodynamic) limit. In this low-energy limit, physics is described by collective variables and quasiparticles, which form the analogue of the Universe. The effective quantum fields in these effective Universes have zero-point energies. These energies are formally diverging (in the same way as in our Universe) and require the ultraviolet cut-off. However, looking at the atomic (ultraviolet) physics we can see that all these divergencies are formal and have no physical consequences. They do not determine the energy of the ground state of the quantum liquid – the energy of the ”quantum vacuum”, which is zero in equilibrium.
@@
One can see that the thermodynamic potenial $\epsilon-\mu n$ plays the role of the energy density of the vacuum, $\epsilon_{\rm vac}$, which obeys the equation of state $w=-1$:
@@ VIII.2 Cosmological constant is zero in equilibrium
Let us first consider the quantum vacuum in the absence of gravity, i.e. in special relativity.Volovik2003c In this case if the external pressure acting on the quantum vacuum is absent, then according to Eq.(69), the cosmological constant is naturally zero. So, the thermodynamics naturally solves the main cosmological constant problem without any fine-tuning, if gravity is absent. Note that the quantity $\epsilon(q)$ is determined by the UV physics, and is huge. But in equilibrium it is naturally cancelled by the term $-\mu q$, which is the analogue of the counterterm in quantum field theories. The mechanism of cancellation is purely thermodynamic and does not depend on whether the vacuum is relativistic or not.
@@ VIII.3 Vacuum decay without fine-tuning
Now consider the influence of gravitational degrees of freedom. As shown in section IV, gravity with its thermodynamically conjugate variables (the gravitational coupling $K$ and the scalar Riemann curvature ${\cal R}$) adds its contribution to the total pressure. Then at equilibrium the total pressure is zero according to eq. (28). But now the equilibrium state with zero pressure corresponds to the de Sitter vacuum, which by virtue of its symmetry serves as an attractor along with the Minkowski vacuum.
The de Sitter attractor can be considered using the dynamics of the Hawking $q$-field. Let us first assume that the Big Bang has started in the originally equilibrium Minkowski vacuum. This corresponds to the special value $\mu_{0}$ of the chemical potential of the $q$ field in Eq.(69) for which $\Lambda(\mu_{0})=0$. Then from our equations without dissipationKlinkhamerVolovik2008b it follows that the cosmological constant, which is very large immediately after the Big Bang, relaxes with oscillations to zero value. Its magnitude averaged over fast oscillations reaches the present value in the present time $\Lambda(t_{\rm present})\right>\sim M^{2}_{\rm P}/t^{2}_{\rm present}\sim 10^{-120}M^{4}_{\rm P}$. The similar behaviour was discussed in Sec. V.3. The oscillating decay also present in the Starobinsky inflation,StarobinskyYokoyama1994 ; Kofman1997 where the magnitude of the oscillation frequency is determined by the Higgs inflaton mass, instead of the Planck scale in our case.
@@ IX.1 Composite tetrads from relative symmetry breaking
This order parameter breaks the separate symmetries under orbital and spin transformations, but remains invariant under the combined rotations. On the level of the Lorentz symmetries the symmetry breaking scheme is $L_{L}\times L_{S}\rightarrow L_{J}$. Here $L_{L}$ is the group of Lorentz transformations in the coordinates space, $L_{S}$ is the group of Lorentz transformations in the spin space, and $L_{J}$ is the symmetry group of the order parameter, which is invariant under the combined Lorentz transformations $L_{J}$.
@@ IX.2 Dimensionful metric and dimensionless interval
Ii is important that in this quantum gravity, the fermionic fields $\Psi$ are dimensionlessVladimirovDiakonov2012 (we denote this as $[\Psi]=[1]$, i.e. $\Psi$ has the same zero dimension as numbers). Thus the tetrads in Eq.(73) have the dimensions of the inverse time and inverse length, $[e^{a}_{0}]=1/[t]$ and $[e^{a}_{i}]=1/[L]$, while the metric elements in Eq.(74) have dimensions $1/[t]^{2}$, $1/[L]^{2}$ and $1/[t][L]$. Due to such dimensions of tetrads and metric, the interval is dimensionless:
@@ IX.3 Dimensions of gauge field and mass
The $U(1)$ field is the geometric quantity, which comes from the gauging of the global $U(1)$ field. The field $A_{\mu}$ comes from the gauging of gradient of the phase field, and thus has dimension of the gradient of phase, with $[A_{0}]=1/[t]$ and $[A_{i}]=1/[L]$. The charge $q$ here is dimensionless – it is the integer (or fractional) geometric charge of the fermionic or bosonic field. In case of electromagnetic $U(1)$-field, $q$ is expressed in terms of the electric charge of electron, i.e. $q=-1$ for electron and $q=+1$ for proton. As a result the action (76) is naturally dimensionless, $[S]=[1]$.
@@
[…]
Since both the interval $ds$ and the action $S$ are dimensionless, from equation (77) it follows that the particle mass $M$ is also dimensionless, $[M]=[S]=[s]=[L]^{0}=[t]^{0}=[1]$. Note that we discuss here the space and time dimensions of the physical quantities. In quantum field theories, the mass dimensions is also used, which will be discussed in Sec. IX.5.
@@ IX.4 Two Planck constants
Comparing the gradient term and the mass terms, and using the dimension of the metric, one again obtains that the mass $M$ is dimensionless, $[M]=[1]$. Then since the action $S$ and volume element $d^{4}x\,\sqrt{-g}$ are dimensionless, it follows that the scalar field is also dimensionless, $[\Phi]^{2}=[M]=[S]=[1]$.
@@
This shows that the Minkowski elements of metric play the role of the Planck constants $\hbar$ and $/\!\!\!h\equiv\hbar c$, which connect the energy and momentum of particles with the frequency and the wave vector. This suggests, that we have the following relations between the Planck constants and the Minkowski metric:
$-g^{00}_{\rm Mink}\equiv\hbar^{2}\,\,,\,\,g^{ik}_{\rm Mink}\equiv/\!\!\!h^{2}\delta^{ik}\,.$ (83)
@@
$c^{2}=\frac{/\!\!\!h^{2}}{\hbar^{2}}\,.$ (84)
[…]
The diffeomorphism invariant quantities are dimensionless. This includes the action $S$ (example is in Eq.(72)); the scalar curvature ${\cal R}$; the scalar field $\Phi$; the wave function $\psi$; masses $M$; the cosmological constant $\Lambda$; the gravitational coupling $K$; etc.Volovik2021 ; Volovik2021cont Here the term ”dimensionless” means the absence of space or time dimensions. However, these dimensionless quantities may have nonzero mass dimension $[M]$. For example, action has zero mass dimension; the mass dimension of curvature is $-2$, i.e. $[{\cal R}]=[M]^{-2}$; the mass dimension of gravitational coupling is 2, i.e. $[K]=[M]^{2}$; the cosmological constant $\Lambda$ has zero mass dimension.
[…]
@@ IX.5 Planck constants do not enter diffeomorphism invariant expressions
$S=K\int d^{4}x\,\sqrt{-g}\,{\cal R}\,\,,\,\,K=\frac{1}{16\pi}\frac{/\!\!\!h}{{\bar{G}}}\equiv\frac{M_{\rm P}^{2}}{16\pi}\,.$ (86)
As we already mentioned, the gravitational coupling $K$ and curvature ${\cal R}$ are space-time dimensionless, but have mass dimensions 2 and $-2$ correspondingly. That is why, although $/\!\!\!h$ and ${\bar{G}}$ have the same space-time dimension, their ratio has mass dimension 2. The length of the Planck constant $/\!\!\!h$ and the Planck length $l_{\rm P}=/\!\!\!h/M_{\rm P}$ coincide if we choose the Planck $M_{\rm P}$ as unity of mass.
@@
$E_{H}=\epsilon_{\rm vac}V_{H}=\frac{1}{2\hbar H}\frac{/\!\!\!h}{\bar{G}}=\frac{M_{\rm P}}{2\pi T}\,M_{\rm P}\equiv NM_{\rm P}\,.$ (89)
@@
$S_{H}=\frac{A}{4/\!\!\!h\bar{G}}=\frac{\pi}{(\hbar H)^{2}}\frac{/\!\!\!h}{\bar{G}}=\frac{M_{\rm P}^{2}}{\pi T^{2}}=4\pi N^{2}\,.$ (90)
The entropy of the Hubble volume is equivalent to the entropy of the cosmological horizon. Quantization of the horizon entropy is considered in many theories. In the Bekenstein approach, Bekenstein1974 ; Mukhanov1986 ; Kastrup1997 ; Khriplovich2008 ; Dvali2011 ; Kiefer2020 ; Bagchi2024 the entropy is the adiabatic invariant, which spectrum is equally spaced giving rise to the linear law, $S_{\rm BH}=aN$. Here $a$ is the dimensionless parameter, which value depends on the microscopic theory. The quantization inspired by the string theoryGibbons2011 ; Visser2012 gives the square-root rule, $S=2\pi(\sqrt{N_{1}}\pm\sqrt{N_{2}})$.
Note that the equation (90) applies to the entropy of the Hubble volume, whereas in general the entropy of the de Sitter state is extensive, see section VI.5.
@@ IX.7 Black hole as ensemble of Planck ”atoms”
The ensemble of the $N$ Planck ”atoms” in Eq.(89) is similar to the ensemble of $N$ Planck size black holes inside the black hole horizon,Volovik2025 where the entropy is also proportional to $N^{2}$:
$S(N)=N^{2}\,.$ (91)
The probability of the quantum tunneling process of the emission of the black hole quantum ($N=1$) is, on the one hand, similar to Eq.(40) for quantum tunneling process of the vortex instanton, and on the other hand, it satisfies the equation (45) for quantum tunneling as random fluctuation:
$w\sim e^{-2N}\approx e^{S(N-1)+S(1)-S(N)}\,.$ (92)
The black hole entropy is non-extensive and thus requires the generalized statistics. The entropy in Eq.(91) satisfies the composition rule of the Tsallis-Cirto $\delta=2$ entropy for the ensembles with $N_{1}$ and $N_{2}$ quanta:
$\sqrt{S_{\delta=2}(N_{1}+N_{2})}=N_{1}+N_{2}=\sqrt{S_{\delta=2}(N_{1})}+\sqrt{S_{\delta=2}(N_{2})}\,.$ (93)
This reproduces the Eq(47). In rare quantum fluctuation processes leading to black hole fragmentation, $N\rightarrow N_{1}+N_{2}$, the entropy decreases sharply. Then the entropy continuously increases in the further process of black hole recombination, $N_{1}+N_{2}\rightarrow N$, reaching a maximum value for a given $N$. In this description, a black hole is not a quantum state, but a dissipative process in which continuous evolution alternates with random quantum jumps, see e.g. Refs.Cohen-Tannoudji1993 ; Frohlich2024 .

### v10

@@ VI.6 Entropy of black hole singularity
[…]
@@
The same singular entropy density in Eq.(52) is obtained for the charged Reissner-Nordström (RN) black hole. According to Ref.Volovik2022G the entropy and the Bekenstein-Hawking temperature of the RN black hole are the same as for the electrically neutral black hole with the same mass. The extra term is to be added to the right hand side of Eq.(50), $\Phi q({\bf r})$, which comes from charge density $q({\bf r})=Q\delta({\bf r})$. However this term is zero, because the corresponding electric analogue of the chemical potential (the electrostatic potential) is zero, $\Phi=0$.
So, if our assumption is correct, then it follows from Eqs. (52) and (53) that all the entropy of the black hole, $S_{\rm BH}=A/4G$, is concentrated in the black hole singularity. This is quite natural from the condensed matter point of view: a singularity is a compact physical object with an enormous energy density. The condensed matter analogue of such object is represented by the core of a vortex in a superfluid liquid. The vortex core contains the singularity in the superfluid velocity, $\nabla\times{\bf v}_{s}=\kappa\hat{\bf z}\delta({\bf r}_{\perp})$, where $\kappa$ is the quantum of circulation of the superfluid velocity. In the chiral Weyl superfluids, the vortex core has a large density of states, which is caused by the flat band in the spectrum of electrons confined in the singularity.Volovik2011
It would be interesting to apply this approach to the Kerr black hole, and to the super-critical Reissner-Nordström black hole with naked singularity.Tumulka2025 ; Volovik2025P
@@ VII.1 Schwinger vs Unruh
[…]
@@ VII.2 Double Unruh temperature from Schwinger+Unruh
To connect the Schwinger pair production with the Unruh effect, let us consider the following combined process in which the particles with charges $\pm q$ and masses $M+m$ are created. At first step the particles with charges $\pm q$ and masses $M$ are created, at the second step these charged particles are accelerated by electric field. These particles play the role of the detectors which energy levels are excited from the levels with energy $M$ to the levels with energy $M+m$ due to the Unruh effect. Since these processes occur in unison, the probability of the combined process $\Gamma^{\rm Schw}(M+m)$ is equal to the product of partial probabilities:
@@ VIII.1 Vacuum energy in condensed matter
In condensed matter systems, such as superfluid 4He and superfluid phases of 3He, we know physics both at the ultraviolet (atomic) scale and in the infrared (hydrodynamic) limit. In this low-energy limit, physics is described by collective variables and quasiparticles, which form the analogue of the Universe. The effective quantum fields in these effective Universes have zero-point energies. These energies are formally diverging (in the same way as in our Universe) and require the ultraviolet cut-off. However, looking from the microscopic (atomic ultraviolet) side we can see that all these divergencies are formal and have no physical consequences. They do not determine the energy of the ground state of the quantum liquid – the energy of the ”quantum vacuum”, which is zero in equilibrium.
@@
One can see that the thermodynamic potenial $\epsilon-\mu n$ plays the role of the energy density of the vacuum, $\epsilon_{\rm vac}$, which obeys the equation of state w=1w=-1:
@@ VIII.2 Cosmological constant is zero in equilibrium
Let us first consider the quantum vacuum in the absence of gravity, i.e. in special relativity.Volovik2003c In this case if the external pressure acting on the quantum vacuum is absent, then according to Eq.(69), the cosmological constant is naturally zero. So, the thermodynamics naturally solves the main cosmological constant problem without any fine-tuning, if gravity is absent. Note that the quantity $\epsilon(q)$ is determined by the UV physics, and is huge. But in equilibrium it is naturally cancelled by the term $\mu q$, which is the analogue of the counterterm in quantum field theories. The mechanism of cancellation is purely thermodynamic and does not depend on whether the vacuum is relativistic or not.
@@ VIII.3 Vacuum decay without fine-tuning
[…]
The de Sitter attractor can be considered using the dynamics of the Hawking $q$-field. Let us first assume that the Big Bang has started in the originally equilibrium Minkowski vacuum. This corresponds to the special value $\mu_{0}$ of the chemical potential of the $q$ field in Eq.(69) for which $\Lambda(\mu_{0})=0$. Then from our equations without dissipationKlinkhamerVolovik2008b it follows that the cosmological constant, which is very large immediately after the Big Bang, relaxes with oscillations to zero value. Its magnitude averaged over fast oscillations reaches the present value in the present time $\left<\Lambda(t_{\rm present})\right>\sim M^{2}_{\rm P}/t^{2}_{\rm present}\sim 10^{-120}M^{4}_{\rm P}$. The similar behaviour was discussed in Sec. V.3. The oscillating decay also present in the Starobinsky inflation,StarobinskyYokoyama1994 ; Kofman1997 where the magnitude of the oscillation frequency is determined by the Higgs inflaton mass, instead of the Planck scale in our case.
@@ IX.1 Composite tetrads from relative symmetry breaking
This order parameter breaks the separate symmetries under orbital and spin transformations, but remains invariant under the combined rotations. On the level of the Lorentz symmetries the symmetry breaking scheme is $L_{L}\times L_{S}\rightarrow L_{J}$. Here $L_{L}$ is the group of Lorentz transformations in the coordinates space, $L_{S}$ is the group of Lorentz transformations in the spin space, and $L_{J}$ is the symmetry group of the order parameter $e^{a}_{\mu}$, which is invariant under the combined Lorentz transformations $L_{J}$.
@@ IX.2 Dimensionful metric and dimensionless interval
Ii is important that in this quantum gravity, the fermionic fields $\Psi$ are dimensionlessVladimirovDiakonov2012 (we denote this as $\Psi]=[1]$, i.e. $\Psi$ has the same zero dimension as numbers). Thus the tetrads in Eq.(73) have the dimensions of the inverse time and inverse length, $e^{a}_{0}]=1/[t]$ and $e^{a}_{i}]=1/[L]$, while the metric elements in Eq.(74) have dimensions $1/[t]^{2}$, $1/[L]^{2}$ and $1/[t][L]$. Due to such dimensions of tetrads and metric, the interval is dimensionless:
@@ IX.3 Dimensions of gauge field and mass
The $U(1)$ field is the geometric quantity, which comes from the gauging of the global $U(1)$ field. The field $A_{\mu}$ comes from the gauging of gradient of the phase field, and thus has dimension of the gradient of phase, with $A_{0}]=1/[t]$ and $A_{i}]=1/[L]$. The charge $q$ here is dimensionless – it is the integer (or fractional) geometric charge of the fermionic or bosonic field. In case of electromagnetic $U(1)$-field, $q$ is expressed in terms of the electric charge of electron, i.e. q=1q=-1 for electron and $q=+1$ for proton. As a result the action (76) is naturally dimensionless, $[S]=[1]$.
@@
$S=M\int ds\,.$ (77)
Since both the interval $ds$ and the action $S$ are dimensionless, from equation (77) it follows that the particle mass $M$ is also dimensionless, $M]=[S]=[s]=[L]^{0}=[t]^{0}=[1]$. Note that we discuss here the space and time dimensions of the physical quantities. In quantum field theories, the mass dimensions is also used, which will be discussed in Sec. IX.5.
@@ IX.4 Two Planck constants
Comparing the gradient term and the mass terms, and using the dimension of the metric, one again obtains that the mass $M$ is dimensionless, $[M]=[1]$. Then since the action $S$ and volume element $d^{4}x\,\sqrt{-g}$ are dimensionless, it follows that the scalar field is also dimensionless, $\Phi]^{2}=[M]=[S]=[1]$.
@@
This shows that the Minkowski elements of metric play the role of the Planck constants $\hbar$ and $h\equiv\hbar c$, which connect the energy and momentum of particles with the frequency and the wave vector. This suggests, that we have the following relations between the Planck constants and the Minkowski metric:
$g^{00}_{\rm Mink}\equiv\hbar^{2}\,\,,\,\,g^{ik}_{\rm Mink}\equiv/\!\!\!h^{2}\delta^{ik}\,.$ (83)
Both Planck constants, bar $\hbar$ and slash /h/\!\!\!h, are the elements of the metric and tetrads in Minkowski vacuum. The bar $\hbar$ is the time component of the inverse tetrad and has dimension of time, $\hbar]=[t]$ while the slash /h/\!\!\!h enters the space components of tetrads and has dimension of length, [/h]=[L][/\!\!\!h]=[L]. The presence of the Planck constants in the metric elements demonstrates that the metric field describes the dynamics of the quantum vacuum, rather than the space-time geometry. This is natural, since space-time itself with its metric arises from the dynamics of quantum fields. The speed of light, as an element of geometry, arises as the ratio of two Planck constants:
$c^{2}=\frac{/\!\!\!h^{2}}{\hbar^{2}}\,.$ (84)
In principle, there can be more than two Planck constants. For example, two Planck constants $\hbar$ and $\hbar_{g}$ were introduced in Ref. Terno2024 for matter and gravity correspondingly (see also Caro1999 ). This suggests that gravity will be more classical than matter if $\hbar_{g}\ll\hbar$. With slash component $h_{g}$ added one obtains four different Planck constants: $\hbar$, /h/\!\!\!h, $\hbar_{g}$ and slash $h_{g}$. Then in principle one may have two speeds of light: $c=/\!\!\!h/\hbar$ and $c_{g}=/\!\!\!h_{g}/\hbar_{g}$, although $c_{g}=c$ is more natural.
@@ IX.5 Planck constants do not enter diffeomorphism invariant expressions
Since $M$ is spacetime dimensionless, from equation $M=\hbar\omega$ it follows that the Planck constant has dimension of time, $\hbar]=[t]$, and the slash Planck constant has dimension of length, [/h]=[L][/\!\!\!h]=[L]. Since these ”constants” are not diffeomorphism invariant, they cannot enter the diffeomorphism invariant expressions. Example is the original action (72), which does not contain the Planck constants $\hbar$ and /h/\!\!\!h. This is the property of any action, if it is written in the diffeomorphism-invariant form. Let us consider several examples, when the Planck constants are removed from the expressions rewritten in diffeomorphism-invariant form.
@@
Since the potential $U$ has dimension of mass $M$, and thus is the space-time dimensionless, one obtains that $\bar{G}$} has the same dimension of length as /h/\!\!\!h, i.e. ${\bar{G}}]=[{/\!\!\!h}]=[L]$. This demonstrates that both $\bar{G}$} and /h/\!\!\!h are not diffeomorphism invariant, and thus they cannot be the fundamental constants. In the gravitational action they compensate each other:
$S=K\int d^{4}x\,\sqrt{-g}\,{\cal R}\,\,,\,\,K=\frac{1}{16\pi}\frac{/\!\!\!h}{{\bar{G}}}\equiv\frac{M_{\rm P}^{2}}{2}\,.$ (86)
@@
$E_{H}=\epsilon_{\rm vac}V_{H}=\frac{1}{2\hbar H}\frac{/\!\!\!h}{\bar{G}}=\frac{4M_{\rm P}}{T}\,M_{\rm P}\equiv NM_{\rm P}\,.$ (89)
@@
$S_{H}=\frac{A}{4/\!\!\!h\bar{G}}=\frac{\pi}{(\hbar H)^{2}}\frac{/\!\!\!h}{\bar{G}}=8\frac{M_{\rm P}^{2}}{T^{2}}=\frac{1}{2}N^{2}\,.$ (90)
This leads to quantization of the cosmological constant.Volovik2025P
The entropy of the Hubble volume is equivalent to the entropy of the cosmological horizon. Quantization of the horizon entropy is considered in many theories. In the Bekenstein approach, Bekenstein1974 ; Mukhanov1986 ; Kastrup1997 ; Khriplovich2008 ; Dvali2011 ; Kiefer2020 ; Bagchi2024 the entropy is the adiabatic invariant, which spectrum is equally spaced giving rise to the linear law, $S_{\rm BH}=aN$. Here $a$ is the dimensionless parameter, which value depends on the microscopic theory. The quantization inspired by the string theoryGibbons2011 ; Visser2012 gives the square-root rule, $S=2\pi(\sqrt{N_{1}}\pm\sqrt{N_{2}})$ Note that the equation (90) applies to the entropy of the Hubble volume, whereas in general the entropy of the de Sitter state is extensive, see section VI.5.
@@ IX.7 Black hole as ensemble of Planck ”atoms”
The ensemble of the $N$ Planck ”atoms” in Eq.(89) is similar to the ensemble of $N$ Planck size black holes inside the black hole horizon,Volovik2025 ; Volovik2025P where the entropy is also proportional to $N^{2}$ at large $N$:
$S(N)=\frac{N(N-1)}{2}\,.$ (91)
The probability of the quantum tunneling process of the emission of the black hole quantum ($N=1$ and $S(N=1)=0$) is, on the one hand, similar to Eq.(40) for quantum tunneling process of the vortex instanton, and on the other hand, it satisfies the equation (45) for quantum tunneling as random fluctuation:
$w\sim e^{S(N-1)-S(N)}=e^{-(N-1)}=\exp\left(-\frac{m_{\rm P}}{T_{\rm H}}\left(1-\frac{m_{\rm P}}{M}\right)\right)\,.$ (92)
The correction to the Hawking radiation rate corresponds to the effect of back reaction in the process of radiation considered by Parikh and Wilczek.ParikhWilczek2000
The black hole entropy is non-extensive and thus requires the generalized statistics. For large $N$, the entropy in Eq.(91) satisfies the composition rule of the Tsallis-Cirto $\delta=2$ entropy for the ensembles with $N_{1}$ and $N_{2}$ quanta:
$\sqrt{S_{\delta=2}(N_{1}+N_{2})}=\sqrt{S_{\delta=2}(N_{1})}+\sqrt{S_{\delta=2}(N_{2})}\,$, (93)
which reproduces the Eq(47). In rare quantum fluctuation processes leading to black hole fragmentation, $N=N_{1}+N_{2}$, the entropy decreases sharply. Then the entropy continuously increases in the further process of black hole recombination, $N_{1}+N_{2}\rightarrow N$, reaching a maximum value for a given $N$. In this description, a black hole is not a quantum state, but a dissipative process in which continuous evolution alternates with random quantum jumps, see e.g. Refs.Cohen-Tannoudji1993 ; Frohlich2024 .